A power transmission line is made of copper 1 .80cm in diameter. Find the resistance of one mile of this line.


1
Expert's answer
2021-05-11T17:20:37-0400

Gives

Diameter of copper wire (d)=1.80cm

Cooper wire resistivity"\rho=1.72\times10^{-8}\Omega m"

Radius (r)=0.90cm "=0.90\times10^{-2}meter"

Length (L)=1mile=1609.34meter

Cross section area of copper wire (A)="\u03c0r^2"

"A=3.14\times{0.009}^2"

"A=25.434\times10^{-5}m^2"


Resistance

"R=\frac{\rho l}{A}\rightarrow(1)"

Put value in eqution (1)

"R=\frac{\rho l}{A}"

"R=\frac{1.72\times10^{-8}\times1609.34}{25.434\times10^{-5}}"

"R=0.1088\Omega"